Directions: Write a system of two equations in two unknowns for each problem. Solve each system by the method of your choice.
The probability that San Fransisco plays in the next super bowl is nine times the probability that they do not play in the next super bowl. The probability that San
Fransisco plays in the next super bowl plus the probability that they do not play is 1. What is the probability that San Fransisco plays in the next super bowl?

Answer by stanbon(75887)

The probability that San Fransisco plays in the next super bowl is nine times the probability that they do not play in the next super bowl.
The probability that San Fransisco plays in the next super bowl plus the probability that they do not play is 1.
What is the probability that San Fransisco plays in the next super bowl?
---------------
Let x be the probability they play; Let y be the probability they do not play.
x+y = 1
x = 9y
------------
Substitute to get:
9y+y=1
y=1/10
x=9y= 9/10 (probabiliy they play in the next Super Bowl)
==========
Cheers,
Stan H.